How many moles of oxygen can be burned biologically when 42.26 g of carbon dioxide is produced


1
Expert's answer
2021-04-14T07:06:30-0400

C6H12O6(s) + 6O2(g) → 6CO2(g) + 6H2O(l)

According to the equation, n (O2) = n (CO2).

n (CO2) = m / M

M (CO2) = 44.01 g/mol

n (O2) = n (CO2) = 42.26 / 44.01 = 0.96 mol